# Gaussian sequence model: sample complexities of testing, estimation and LFHT

### v1

@@ Appendix B Analysis of LFHT for General Convex Sets
$\chi^{2}(\mathbb{P}_{0,Z\mid X}\ \|\ \mathbb{P}_{1,Z\mid X}\mid\mathbb{P}_{0,X})+1$
@@
$\mathbb{P}(\theta_{j}=\theta_{j}^{\prime}\mid\mathbf{X})\cdot\exp(m(\theta_{i}^{\star})^{2})+\mathbb{P}(\theta_{j}\neq\theta_{j}^{\prime}\mid\mathbf{X})\cdot\exp(-m(\theta_{i}^{\star})^{2})$
@@ Appendix C Analysis of LFHT for $\ell_{p}$ Bodies
$\sum_{t\in T_{2}}\left((p_{\mathtt{X}})_{t}-(p_{\mathtt{Y}})_{t}\right)^{2}\cdot\mathbb{I}\left[|(p_{\mathtt{X}})_{t}-(p_{\mathtt{Y}})_{t}|<6\sqrt{\frac{2\log(4D/\delta)}{n}}\right]$
@@
According to the construction of set $T$, we have for any $t\in[D]\backslash T$,
@@
$\sum_{t\in T}\left((p_{\mathtt{X}})_{t}-(p_{\mathtt{Y}})_{t}\right)^{2}$
@@
$\chi^{2}(\mathbb{P}_{0,Z\mid X}\ \|\ \mathbb{P}_{1,Z\mid X}\mid\mathbb{P}_{0,X})+1$
@@
$\mathbb{P}(\theta_{j}=1\mid\mathbf{X})$
@@
$\mathbb{P}(\theta_{j}=\theta_{j}^{\prime}\neq 0\mid X)\left(\exp(mr^{2})-1\right)+\mathbb{P}(\theta_{j}=-\theta_{j}^{\prime}\neq 0\mid X)\left(\exp(-mr^{2})-1\right)+1$
@@
$\left(p_{j}(\mathbf{X})+q_{j}(\mathbf{X})\right)^{2}\leq\frac{h^{2}}{4}\cdot\left(\exp\left(r\cdot\sum_{i=1}^{n}X_{j}^{i}\right)+\exp\left(-r\cdot\sum_{i=1}^{n}X_{j}^{i}\right)\right)^{2}$
@@
$\mathbb{E}\left[\mathbb{P}(\theta_{j}=\theta_{j}^{\prime}\neq 0\mid X)\left(\exp(mr^{2})-1\right)+\mathbb{P}(\theta_{j}=-\theta_{j}^{\prime}\neq 0\mid X)\left(\exp(-mr^{2})-1\right)+1\right]$
@@
$\chi^{2}(\mathbb{P}_{0,Z\mid X}\ \|\ \mathbb{P}_{1,Z\mid X}\mid\mathbb{P}_{0,X})]$

### v4

@@ Appendix B Analysis of LFHT for General Convex Sets
$\hskip-14.22636pt\chi^{2}(\mathbb{P}_{0,Z\mid X}\ \|\ \mathbb{P}_{1,Z\mid X}\mid\mathbb{P}_{0,X})+1$
@@
$\hskip-14.22636pt\mathbb{P}(\theta_{j}=\theta_{j}^{\prime}\mid\mathbf{X})\cdot\exp(m(\theta_{i}^{\star})^{2})+\mathbb{P}(\theta_{j}\neq\theta_{j}^{\prime}\mid\mathbf{X})\cdot\exp(-m(\theta_{i}^{\star})^{2})$
@@ Appendix C Analysis of LFHT for $\ell_{p}$ Bodies
$\hskip-14.22636pt\sum_{t\in T_{2}}\left((p_{\mathtt{X}})_{t}-(p_{\mathtt{Y}})_{t}\right)^{2}\cdot\mathbb{I}\left[|(p_{\mathtt{X}})_{t}-(p_{\mathtt{Y}})_{t}|<6\sqrt{\frac{2\log(4D/\delta)}{n}}\right]$
@@
According to the construction of set $T$, we have for any t[$\Tt\in[D]\backslash$ T,
@@
$\hskip-14.22636pt\sum_{t\in T}\left((p_{\mathtt{X}})_{t}-(p_{\mathtt{Y}})_{t}\right)^{2}$
@@
$\hskip-14.22636pt\chi^{2}(\mathbb{P}_{0,Z\mid X}\ \|\ \mathbb{P}_{1,Z\mid X}\mid\mathbb{P}_{0,X})+1$
@@
$\hskip-14.22636pt\mathbb{P}(\theta_{j}=1\mid\mathbf{X})$
@@
$\hskip-14.22636pt\mathbb{P}(\theta_{j}=\theta_{j}^{\prime}\neq 0\mid X)\left(\exp(mr^{2})-1\right)+\mathbb{P}(\theta_{j}=-\theta_{j}^{\prime}\neq 0\mid X)\left(\exp(-mr^{2})-1\right)+1$
@@
$\hskip-14.22636pt\left(p_{j}(\mathbf{X})+q_{j}(\mathbf{X})\right)^{2}\leq\frac{h^{2}}{4}\cdot\left(\exp\left(r\cdot\sum_{i=1}^{n}X_{j}^{i}\right)+\exp\left(-r\cdot\sum_{i=1}^{n}X_{j}^{i}\right)\right)^{2}$
@@
$\hskip-14.22636pt\mathbb{E}\left[\mathbb{P}(\theta_{j}=\theta_{j}^{\prime}\neq 0\mid X)\left(\exp(mr^{2})-1\right)+\mathbb{P}(\theta_{j}=-\theta_{j}^{\prime}\neq 0\mid X)\left(\exp(-mr^{2})-1\right)+1\right]$
@@
$\hskip-14.22636pt\chi^{2}(\mathbb{P}_{0,Z\mid X}\ \|\ \mathbb{P}_{1,Z\mid X}\mid\mathbb{P}_{0,X})]$